A stock had returns of 17.39 percent, −11.26 percent, 22.94 percent, and 13.59 percent for the past four years. What is the standard deviation of the returns? Multiple Choice 13.60% 2.28% 12.09% 15.11% 22.84%
Answer #1

Average return=Total return/Total time period

=(17.39-11.26+22.94+13.59)/4=10.665%

Return (Return-Average return)^2
17.39 (17.39-10.665)^2=45.225625
-11.26 (-11.26-10.665)^2=480.705625
22.94 (22.94-10.665)^2=150.675625
13.59 (13.59-10.665)^2=8.555625
Total=685.1625%

Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[685.1625/(4-1)]^(1/2)

=[685.1625/3]^(1/2)

=15.11%(Approx)